Chapter 13, Problem 018 Flying Circus of Physics Mountain pull. A large mountain can slightly affect the direction of down as determined by a plumb line. Assume that we can model a mountain as a sphere of radius R 1.00 km and density (mass per unit volume) 2.6 × 103 kg/m3. Assume also that we hang a 0.750m plumb line at a distance of 3R from the spheres center and such that the sphere pulls horizontally on the lower end. How far would the lower end move toward the sphere? Number Units the tolerance is +/-2% SHOW HINT LINK TO TEXT LINK TO SAMPLE PROBLEM MATH HELP Question Attempts: 0 of 5 used SAVE FOR LATER SUBMIT ANSWER
Answer #1

T cos θ Fg 30 Tsin θ

Density of mountain -2.6 * 103 kg/m

Radius of mountain R1.00km 103 m

Mass of mountain M= ho V= ho*rac{4}{3}pi R^3

The sphere (mountain) pulls the plumb line,

Forces acting the plumb mass m are,

i) gravitational force tg , vertically downwards

ii) Components of tension Tcos heta vertically upwards, Tsin heta horizontally towards right.

iii) Force of attraction due to mountain (sphere) GMm 9(3R) horizontally towards left.

Net force acting on the plumb bob is zero, F_g=Tsin heta and ng= T cos θ

Tsino F

an heta=rac{GMm}{mg(3R)^2}

an heta=rac{GM}{9gR^2}

27g

an heta=rac{4pi *6.67*10^{-11}*2.6*10^3*10^3}{27*9.81}=8.23*10^{-6}

From the figure, an heta=rac{x}{l} ( where l is length of plumb line, x is the distance moved by lower end of plumb line towards sphere )

2.-1 tan θ-0.750 * 8.23 * 10-6-6.17 * 10-6 m